show me how to find answer to:3 4/7 divided by 3 1/2. is it...3 4/7 divided by 3 1/2-25/7 x 2/7=50/49 which is 1 1/49?
Answer by Alan3354(69443)

show me how to find answer to:
3 4/7 divided by 3 1/2. is it...3 4/7 divided by 3 1/2-25/7 x 2/7=50/49 which is 1 1/49?
--------------
3 4/7 divided by 3 1/2
Change to fraction, not mixed numbers
(25/7) / (7/2)
Invert the divisor and multiply
(25/7)*(2/7)
= 50/49
---------
So, yes.